OpenStudy (anonymous):

will medal and fan what is the solution to the proportion 4/12 =10/x A. 3 B. 4.8 C. 3.3 D. 30 @chosenmatt @loganisawesome151

OpenStudy (anonymous):

\[\frac{1}{3}=\frac{10}{x}\\ x=10\times 3\]

OpenStudy (horsegirl27):

10*3 is 30 so your answer will be D

OpenStudy (anonymous):

10*3= 30

OpenStudy (horsegirl27):

Yes

OpenStudy (anonymous):

BOOM didnt even have to try :)

OpenStudy (anonymous):

thx @horsegirl27 and @satellite73

OpenStudy (anonymous):

lol @loganisawesome151

OpenStudy (anonymous):

one more pls??

OpenStudy (horsegirl27):

Your welcome. And to explain more off what @satellite73 said, you simplify 4/12 to 1/3 and you mulitply it to 10/x. Then you get 10*3, which of course is 30

OpenStudy (horsegirl27):

@pielover123 post as a different question

OpenStudy (anonymous):

ok

OpenStudy (horsegirl27):

Just tag me and I can help